# Pressure and speed of sound in two-flavor color-superconducting quark matter at next-to-leading order

### v1

@@ I Introduction and Summary
Let us conclude by commenting on how the inclusion of strange quarks may affect our present results. In three-flavor quark matter, the expansion of the pressure in powers of the gap in Eq. (1) should assume the same functional form as considered in our present work [54, 34], provided that quark mass corrections are suppressed as they are at large densities [32]. The gap in the expansion would in that case then refer to the CFL gap rather than the 2SC gap. Since the CFL gap in the weak-coupling limit is also of the form given in Eq. (2), it is reasonable to expect that our conclusions with respect to the pressure and the speed of sound remain unaltered on a qualitative level. Note that, at the order considered in this work, the coefficient $\gamma_{0}^{\text{3-flavor}}$ for three-flavor quark matter agrees identically with the one for two quark flavors [69, 70, 71, 72, 73, 74] as the flavor dependence is absorbed in $p_{\text{free}}^{\text{3-flavor}}=3\mu^{4}/(4\pi^{2})$. The coefficient $\gamma_{1}^{\text{3-flavor}}$ is known at ${\mathcal{O}}(g^{0})$ for quark matter in the CFL phase, $\gamma_{1}^{\text{3-flavor}}=4$ [86, 54, 57, 59]. Thus, gapped quark matter with two and three flavors behaves qualitatively the same at NLO/LO. Of course, a quantitative comparison requires the computation of the corresponding coefficient $\gamma_{1}^{\text{3-flavor}}$ at $\mathcal{O}(g^{2})$ in $2+1$ flavor QCD, which will be performed in an upcoming work.
@@ II.1 General considerations
$p(\mu)=-\lim_{\beta\to\infty}\frac{1}{\beta}\frac{\partial}{\partial V}\ln{\mathcal{Z}}(\mu)\,,$ (10)
@@
$\lambda_{\text{2SC}}(\tilde{\bar{\psi}}_{b}\tau_{2}i\epsilon_{abc}\gamma_{5}{\mathcal{C}}\tilde{\bar{\psi}}^{T}_{c})(\tilde{\psi}^{T}_{d}{\mathcal{C}}\gamma_{5}\tau_{2}i\epsilon_{ade}\tilde{\psi}_{e})\;\mapsto\;m^{2}\tilde{\Delta}^{\ast}_{a}\tilde{\Delta}_{a}+ih\left[(\tilde{\psi}^{T}_{b}{\mathcal{C}}\gamma_{5}\tau_{2}\tilde{\Delta}_{a}\epsilon_{abc}\tilde{\psi}_{c})-(\tilde{\bar{\psi}}_{b}\gamma_{5}\tau_{2}\tilde{\Delta}^{\ast}_{a}\epsilon_{abc}{\mathcal{C}}\tilde{\bar{\psi}}^{T}_{c})\right]\,,$ (16)
@@ II.3 Expansion of the QCD effective action
$-\frac{1}{2}i{h}(\bar{\psi}_{b}\gamma_{5}\tau_{2}\Delta^{\ast}_{a}\epsilon_{abc}{\mathcal{C}}\bar{\psi}^{T}_{c})+S_{\text{gf}}+S_{\text{gh}}\,.$ (30)
@@
Before proceeding to the computation of the corrections, let us comment on the parameter $m^{2}$ in the action $S^{\prime}$ and its relation to the 2SC gap. This parameter together with the quark-diquark interaction $h$ emerges from the Hubbard-Stratonovich transformation of the four-quark channel associated with 2SC pairing. From Eq. (II.3), we now deduce that $m^{2}$ can be associated with the curvature of an effective action in terms of the gauge-invariant quantity $|\Delta|^{2}$. The ground state configuration of the diquark field derived from such an effective action determines the gap in the quark excitation spectrum. In the present work, we do not aim at a self-consistent computation of this minimum or the gap, both of which would determine $m^{2}$ in terms of the gauge coupling but treat the gap as an external (constant) input parameter for simplicity, $|\Delta_{\text{gap}}|\equiv h|\Delta_{\text{gs}}|$. To be more specific, in our computations of the pressure, we shall rather fix $m^{2}$ to provide a gap that agrees with different previous calculations of the quantity as a function of $\mu$ and $g$. In principle, the self-consistently computed solution of the quantum equations of motion of the diquark field is in general not constant. Therefore, our assumption of a constant gap is a simplification which may potentially affect our results for $\gamma_{1}$ and may introduce a residual gauge dependence in this coefficient. We shall at least investigate the gauge dependence further in an upcoming work on three-flavor quark matter.
@@ III Perturbative computation of the pressure of gapped quark matter
The starting point of our perturbative computation of the pressure of gapped two-flavor quark matter up to ${\mathcal{O}}(g^{2})$ in the strong coupling and $|\Delta_{\text{gap}}|^{2}$ in the gap is given by the action $S^{\prime}$ in Eq. (II.3). For such a perturbative study, it is convenient to introduce Feynman rules as usual. However, since we have introduced the auxiliary background field $\Delta_{a}$ in the action, the quark propagator now assumes the following form:
@@
determines the parameter $m^{2}$ in Eq. (II.3) as a function of $|\Delta_{\text{gap}}|^{2}$ at this order. Recall that we shall treat the gap as an external parameter to fix $m^{2}$ in our present study, as discussed in Subsec. II.3. In any case, the pressure is then obtained as a function of the dimensionless parameter $|\bar{\Delta}_{\text{gap}}|^{2}\equiv|\Delta_{\text{gap}}|^{2}/\mu^{2}$ by evaluating the effective potential at the non-trivial ground state, see Eq. (20), i.e., $p=-U(|\Delta_{\text{gap}}|^{2}/h^{2})$. Up to order $|\bar{\Delta}_{\text{gap}}|^{2}$, we recover
@@
Here, thick lines represent the propagators with the full background-field dependence, while thin lines correspond to background-field independent propagators and thin lines labeled with $\Delta$ are associated with the “propagators” ${\mathcal{P}}^{\Delta}_{\psi}$. The first diagram on the right-hand side of Eq. (LABEL:eq:2loopdiagamsapprox) does not carry a dependence on the background field and corresponds to the standard two-loop contribution. From the standpoint of the effective potential, this contribution induces a shift of the potential of ${\mathcal{O}}(g^{2})$, which in turn yields a correction of ${\mathcal{O}}(g^{2})$ to the coefficient $\gamma_{0}$ in Eq. (1). The second term on the right-hand side of Eq. (LABEL:eq:2loopdiagamsapprox) is explicitly proportional to $|\Delta|^{2}$ and generates the correction of ${\mathcal{O}}(g^{2})$ to $\gamma_{1}$.
@@ Appendix A Two-loop Contribution to
In this Appendix, we comment on the calculation of the two-loop effective potential. By applying the approximations discussed in Sec. III, we obtain the following two-loop integral for the leading background-field dependent contribution on the right-hand side of Eq. (LABEL:eq:2loopdiagamsapprox): [^9] The first-term on the right-hand side of Eq. (LABEL:eq:2loopdiagamsapprox) is obtained by replacing $\mathcal{P}_{\psi}^{\Delta}$ with $\mathcal{P}_{\psi}^{0}$ in Eq. (52).
@@
To compute the two-loop integral in Eq. (52), we have employed a sharp three-momentum cutoff and ensured that our results agree with already existing results, e.g., with the standard one-loop quark self-energy correction in the absence of a background field. The latter not only underlies the computation of the two-loop pressure of ungapped quark matter but also enters the computation of Eq. (53), see also Eq. (LABEL:eq:2loopdiagamsapprox). To perform the numerical integration mentioned above, we have first rescaled the dimensionful variables in the integral to be in units of $\mu$. The numerical integration has been performed for several large values of the dimensionless sharp cutoff $\bar{\Lambda}\equiv\Lambda/\mu$ and small values of the dimensionless background field $|\bar{\Delta}|\equiv|\Delta|/\mu$, i.e., the parameter range for which our approximation is suitable. The numerical data has then been fitted by positing a suitable fitting function $f(\bar{\chi},\bar{\Lambda})$, where $\bar{\chi}\equiv h^{2}|\bar{\Delta}|^{2}$. For vanishing background field, the integral vanishes by construction, which we have also verified numerically. In particular, this implies that there is no $\bar{\Lambda}^{4}$ vacuum contribution needed in the fit function.

### v2

@@ I Introduction and Summary
Let us conclude by commenting on how the inclusion of strange quarks may affect our present results. In three-flavor quark matter, the expansion of the pressure in powers of the gap in Eq. (1) should assume the same functional form as considered in our present work [54, 34], provided that quark mass corrections are suppressed as they are at large densities [32]. The gap in the expansion would in that case then refer to the CFL gap rather than the 2SC gap. Since the CFL gap in the weak-coupling limit is also of the form given in Eq. (2), it is reasonable to expect that our conclusions with respect to the pressure and the speed of sound remain unaltered on a qualitative level in the presence of the CFL diquark gap. [^1] We remark here that in CFL matter there are additional condensates arising from the chiral and baryon symmetry breaking in the CFL phase [53, 59], which may complicate the picture. These condensates can however also be handled using the approach employing in the present work, by introducing additional expansion parameters. Note that, at the order considered in this work, the coefficient $\gamma_{0}^{\text{3-flavor}}$ for three-flavor quark matter agrees identically with the one for two quark flavors [69, 70, 71, 72, 73, 74] as the flavor dependence is absorbed in $p_{\text{free}}^{\text{3-flavor}}=3\mu^{4}/(4\pi^{2})$. The coefficient $\gamma_{1}^{\text{3-flavor}}$ is known at ${\mathcal{O}}(g^{0})$ for quark matter in the CFL phase, $\gamma_{1}^{\text{3-flavor}}=4$ [86, 54, 57, 59]. Thus, gapped quark matter with two and three flavors behaves qualitatively the same at NLO/LO. Of course, a quantitative comparison requires the computation of the corresponding coefficient $\gamma_{1}^{\text{3-flavor}}$ at $\mathcal{O}(g^{2})$ in $2+1$ flavor QCD, which will be performed in an upcoming work.
@@ II.1 General considerations
$p(\mu)=\lim_{\beta\to\infty}\frac{1}{\beta}\frac{\partial}{\partial V}\ln{\mathcal{Z}}(\mu)\,,$ (10)
@@
$\lambda_{\text{2SC}}(\tilde{\bar{\psi}}_{b}\tau_{2}i\epsilon_{abc}\gamma_{5}{\mathcal{C}}\tilde{\bar{\psi}}^{T}_{c})(\tilde{\psi}^{T}_{d}{\mathcal{C}}\gamma_{5}\tau_{2}i\epsilon_{ade}\tilde{\psi}_{e})$
$\mapsto m^{2}\tilde{\Delta}^{\ast}_{a}\tilde{\Delta}_{a}+ih\Big{[}(\tilde{\psi}^{T}_{b}{\mathcal{C}}\gamma_{5}\tau_{2}\tilde{\Delta}_{a}\epsilon_{abc}\tilde{\psi}_{c})$
$-(\tilde{\bar{\psi}}_{b}\gamma_{5}\tau_{2}\tilde{\Delta}^{\ast}_{a}\epsilon_{abc}{\mathcal{C}}\tilde{\bar{\psi}}^{T}_{c})\Big{]}\,,$ (16)
@@ II.3 Expansion of the QCD effective action
$-\frac{1}{2}i{h}(\bar{\psi}_{b}\gamma_{5}\tau_{2}\Delta^{\ast}_{a}\epsilon_{abc}{\mathcal{C}}\bar{\psi}^{T}_{c})\Big{\}}+S_{\text{gf}}+S_{\text{gh}}\,.$ (30)
@@
Before proceeding to the computation of the corrections, let us comment on the parameter $m^{2}$ in the action $S^{\prime}$ and its relation to the 2SC gap. This parameter together with the quark-diquark interaction $h$ emerges from the Hubbard-Stratonovich transformation of the four-quark channel associated with 2SC pairing. From Eq. (30), we now deduce that $m^{2}$ can be associated with the curvature of an effective action in terms of the gauge-invariant quantity $|\Delta|^{2}$. The ground state configuration of the diquark field derived from such an effective action determines the gap in the quark excitation spectrum. In the present work, we do not aim at a self-consistent computation of this minimum or the gap, both of which would determine $m^{2}$ in terms of the gauge coupling but treat the gap as an external (constant) input parameter for simplicity, $|\Delta_{\text{gap}}|\equiv h|\Delta_{\text{gs}}|$. To be more specific, in our computations of the pressure, we shall rather fix $m^{2}$ to provide a gap that agrees with different previous calculations of the quantity as a function of $\mu$ and $g$. In principle, the self-consistently computed solution of the quantum equations of motion of the diquark field is in general not constant. Therefore, our assumption of a constant gap is a simplification which may potentially affect our results for $\gamma_{1}$ and may introduce a residual gauge dependence in this coefficient. We shall at least investigate the gauge dependence further in an upcoming work on three-flavor quark matter.
@@ III Perturbative computation of the pressure of gapped quark matter
The starting point of our perturbative computation of the pressure of gapped two-flavor quark matter up to ${\mathcal{O}}(g^{2})$ in the strong coupling and $|\Delta_{\text{gap}}|^{2}$ in the gap is given by the action $S^{\prime}$ in Eq. (30). For such a perturbative study, it is convenient to introduce Feynman rules as usual. However, since we have introduced the auxiliary background field $\Delta_{a}$ in the action, the quark propagator now assumes the following form:
@@
determines the parameter $m^{2}$ in Eq. (30) as a function of $|\Delta_{\text{gap}}|^{2}$ at this order. Recall that we shall treat the gap as an external parameter to fix $m^{2}$ in our present study, as discussed in Subsec. II.3. In any case, the pressure is then obtained as a function of the dimensionless parameter $|\bar{\Delta}_{\text{gap}}|^{2}\equiv|\Delta_{\text{gap}}|^{2}/\mu^{2}$ by evaluating the effective potential at the non-trivial ground state, see Eq. (20), i.e., $p=-U(|\Delta_{\text{gap}}|^{2}/h^{2})$. Up to order $|\bar{\Delta}_{\text{gap}}|^{2}$, we recover
@@
[…]
Here, thick lines represent the propagators with the full background-field dependence, while thin lines correspond to background-field independent propagators and thin lines labeled with $\Delta$ are associated with the “propagators” ${\mathcal{P}}^{\Delta}_{\psi}$. The first diagram on the right-hand side of Eq. (46) does not carry a dependence on the background field and corresponds to the standard two-loop contribution. From the standpoint of the effective potential, this contribution induces a shift of the potential of ${\mathcal{O}}(g^{2})$, which in turn yields a correction of ${\mathcal{O}}(g^{2})$ to the coefficient $\gamma_{0}$ in Eq. (1). The second term on the right-hand side of Eq. (46) is explicitly proportional to $|\Delta|^{2}$ and generates the correction of ${\mathcal{O}}(g^{2})$ to $\gamma_{1}$.
@@ Appendix A Two-loop Contribution to
In this Appendix, we comment on the calculation of the two-loop effective potential. By applying the approximations discussed in Sec. III, we obtain the following two-loop integral for the leading background-field dependent contribution on the right-hand side of Eq. (46): [^10] The first-term on the right-hand side of Eq. (46) is obtained by replacing $\mathcal{P}_{\psi}^{\Delta}$ with $\mathcal{P}_{\psi}^{0}$ in Eq. (52).
@@
To compute the two-loop integral in Eq. (52), we have employed a sharp three-momentum cutoff and ensured that our results agree with already existing results, e.g., with the standard one-loop quark self-energy correction in the absence of a background field. The latter not only underlies the computation of the two-loop pressure of ungapped quark matter but also enters the computation of Eq. (53), see also Eq. (46). To perform the numerical integration mentioned above, we have first rescaled the dimensionful variables in the integral to be in units of $\mu$. The numerical integration has been performed for several large values of the dimensionless sharp cutoff $\bar{\Lambda}\equiv\Lambda/\mu$ and small values of the dimensionless background field $|\bar{\Delta}|\equiv|\Delta|/\mu$, i.e., the parameter range for which our approximation is suitable. The numerical data has then been fitted by positing a suitable fitting function $f(\bar{\chi},\bar{\Lambda})$, where $\bar{\chi}\equiv h^{2}|\bar{\Delta}|^{2}$. For vanishing background field, the integral vanishes by construction, which we have also verified numerically. In particular, this implies that there is no $\bar{\Lambda}^{4}$ vacuum contribution needed in the fit function.